OpenStudy (anonymous):

Simplify completely:

OpenStudy (anonymous):

\[\sqrt{9}-\sqrt{12}+\sqrt{16}x \sqrt{2}\]

OpenStudy (anonymous):

\[a.)\sqrt{3}+4\] \[b.)3-2\sqrt{3}+4\sqrt{2}\] \[c.)\sqrt{3}-4\sqrt{2}\] \[d.)\sqrt{15}+4\]

OpenStudy (anonymous):

^^i got b for my answer is that correct?

OpenStudy (anonymous):

@megangray32 @Hero

hero (hero):

Yes, however, next time, don't use \(x\) to mean \(\times\) because \(x\) represents a variable, not "multiply".

OpenStudy (anonymous):

thank you @Hero